Marty is proving that the following trigonometric identity is true:
tan^2θ⋅cos^2θ=1−cos^2θ

Which step would be the first line of his proof?
tan^2θ⋅cos^2θ=1−sin^2θ
tan^2θ⋅cos^2θ=tan^2θ
tan^2θ⋅cos^2θ=sin^2θ
tan^2θ=1−cos^2θ⋅cos^2θ

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

I would add cos^2 to each side,
--> tan^2 = 1
Not an identity.
==================
And, I would use parentheses.

Answer by ikleyn(52848)

.
Marty is proving that the following trigonometric identity is true:
tan^2θ⋅cos^2θ=1−cos^2θ

Which step would be the first line of his proof?

tan^2θ⋅cos^2θ=1−sin^2θ

tan^2θ⋅cos^2θ=tan^2θ

tan^2θ⋅cos^2θ=sin^2θ     <<<---+++ This one.

tan^2θ=1−cos^2θ⋅cos^2θ